OpenStudy (anonymous):

find the value of y when x equals -8. the problem is -5x-5y=50

OpenStudy (anonymous):

help me solve this?

OpenStudy (anonymous):

Just plug in the given value. -5(-8) - 5y = 50 40 - 5y = 50 Then just use algebra to find y. -5y = 10 y = -2 You can plug it back in to confirm. -5(-8) - 5(-2) = 50 40 - (-10) = 50

OpenStudy (anonymous):

step 1: -5x-5y=50 step 2: -5(-8)-5y=50 - fill in the X with -8 step 3: 40-5y=50 - times -5 and -8 you get 40 step 4: 40-50=5y - take the -5y and make it positive by moving it over the = sign and take the +50 and move it to the other side of the = sign and turn it to -50 step 5: 8-10=y - we just divided 40 and 50 by the 5y. 40/5=8. 50/5=10. step 6: -2=y - just do 8-10 step 7: -5(-8)-5(-2)=? - go back and check your answer by taking the variables and add it to the equation. step 8: -5(-8)-5(-2)=50 - just solve it and see if its right